OpenStudy (anonymous):

What number must you add to complete the square? x^2+10x=14

OpenStudy (amistre64):

10x/2x = 5

OpenStudy (anonymous):

do i write that who thinq that completes the square or just put 5

OpenStudy (amistre64):

the number that you must add to "complete the square" is: 5 how you want to present that information is up to you :)

OpenStudy (amistre64):

make it 25...

OpenStudy (amistre64):

5 is the number you square to get the number to add..... i forgot a step :)

OpenStudy (amistre64):

the number that you must add to "complete the square" is: 25

OpenStudy (anonymous):

so whats tha tht complete tha square 5 or 25

OpenStudy (amistre64):

x^2+10x +25 =14 +25 ^^^ ^^^ just like this

OpenStudy (anonymous):

idk , i dont know how this math qoess tha answer is 25

OpenStudy (amistre64):

yes, the answer is 25 :)